OpenStudy (anonymous):

Simplify: 2x(4x2 - 3x - 9)

OpenStudy (anonymous):

\[a\left(b+c+d\right)=ab+ac+ad\]Now, use this rule to solve your question.

OpenStudy (anonymous):

is the answer 8x3 - 6x2 - 18x ??

OpenStudy (wasiqss):

yes